OpenStudy (anonymous):

what is the solution of the system? use substitutions y=x+6 and y=2x

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

y=x+6 y=(6)+6 y=36 y=2x y=2(6) y=36